OpenStudy (anonymous):

Complete the standard form of the equation –5y = 10 – 2x how would i do this

OpenStudy (gorv):

divide both side by -5

OpenStudy (anonymous):

Hello, The equation you show fits into the slope-intercept. (i.e. \(y = mx + c\)) Ergo the equation is a linear equation. The standard form of a linear equation is \(Ax + By = C\). Do you know how to rewrite the equation into that format? @quentavess

OpenStudy (anonymous):

um not really can you show me how

OpenStudy (mayankdevnani):

do you know how to write a standard form of line equation ?? @quentavess

OpenStudy (anonymous):

no

OpenStudy (mayankdevnani):

the standard form of equation of line :- \[\large \bf \color{red}{y}=\color{blue}{mx}+\color{green}{c}\]

OpenStudy (mayankdevnani):

and your question :- \[\large \bf -5y=10-2x\]

OpenStudy (mayankdevnani):

can you write this equation into standard form ??

OpenStudy (anonymous):

no not really im terrible

OpenStudy (mayankdevnani):

first,divide whole equation by 5,we get \[\large \bf -y=\frac{10-2x}{5}\]

OpenStudy (mayankdevnani):

right ??

OpenStudy (anonymous):

yeah

OpenStudy (mayankdevnani):

after simplifying this,we get \[\large \bf -y=\frac{10}{5}-\frac{2x}{5}\] \[\large \bf -y=2-\frac{2x}{5}\]

OpenStudy (mayankdevnani):

right ?? @quentavess

OpenStudy (anonymous):

right

OpenStudy (mayankdevnani):

so, multiply by -1 to both sides,we get \[\large \bf y=-2+\frac{2x}{5}\]

OpenStudy (mayankdevnani):

just compare with this equation by standard from of line :- \[\large \bf y=mx+c\] \[\large \bf y=\frac{2}{5}x-2\]

OpenStudy (mayankdevnani):

where, \[\large \bf m=\frac{2}{5}~and~c=-2\]

OpenStudy (mayankdevnani):

hope you understand. @quentavess

OpenStudy (anonymous):

sort of yeah

OpenStudy (mayankdevnani):

good !